# Insights into DeepSeek-V3: Scaling Challenges and Reflections on Hardware for AI Architectures

### v1

@@ Abstract.
[…]
The definitive version will appear as part of the Industry Track in Proceedings of the 52nd Annual International Symposium on Computer Architecture (ISCA ’25).
footnotetext: Authors are listed in alphabetical order of their first names. Yuqing Wang and Liyue Zhang are the corresponding authors of this paper (e-mail: research@deepseek.com).
@@ 1.1. Background
Large Language Models (LLMs) have undergone rapid evolution in recent years, driven by iterative advancements in model design, computational power, and data availability. In 2024, groundbreaking models such as GPT4o (OpenAI, 2024a), LLaMa-3 (AI@Meta, 2024a), Claude 3.5 Sonnet (Anthropic, 2024), Grok-2 (xAI, 2024a), Qwen2.5 (Yang et al., 2024), Gemini-2 (Google, 2024) and our DeepSeek-V3 (DeepSeek-AI, 2024d) have showcased remarkable progress, further narrowing the gap towards Artificial General Intelligence (AGI). As the Scaling Laws (Kaplan et al., 2020) shows, increasing model size, training data, and computational resources leads to substantial improvements in model performance, underscoring the pivotal role of scaling in advancing AI capabilities. Collectively, these developments have ushered in an era where scaling model size and computational power is seen as the key to unlocking higher levels of intelligence.
@@
Recent developments, reasoning models such as OpenAI’s o1/o3 series models (OpenAI, 2024b, 2025), DeepSeek-R1 (DeepSeek-AI, 2025a), Claude-3.7 Sonnet (Anthropic, 2025), Gemini 2.5 Pro (Google, 2025), Seed1.5-Thinking (Seed, 2025) and Qwen3 (Team, 2025) have demonstrated not only the benefits conferred by large-scale architectures, but also the necessity of improving inference efficiency, particularly in handling longer contexts and achieving greater reasoning depth. These advancements underscore the need for faster and more efficient inference, consequently placing ever-increasing demands on computational resources.
@@
To meet these challenges, industry leaders such as Alibaba, ByteDance, Google, xAI and Meta have deployed colossal training clusters (Jouppi et al., 2023; Mudigere et al., 2023; Gangidi et al., 2024; Jiang et al., 2024; Qian et al., 2024; xAI, 2024b), featuring tens or even hundreds of thousands of GPUs or TPUs. While such massive infrastructures have enabled the development of state-of-the-art models, their exorbitant costs present significant barriers for smaller research teams and organizations. Despite these barriers, open-source startups such as DeepSeek (DeepSeek-AI, 2024b, c, d, a, 2025a) and Mistral (Jiang et al., 2023; Mistral, 2024) are also striving to develop state-of-the-art models. Among them, DeepSeek has especially demonstrated that effective software-hardware co-design can enable cost-efficient training of large models, leveling the playing field for smaller teams.
@@
Building on this tradition, DeepSeek-V3 (DeepSeek-AI, 2024d) represents a new milestone in cost-effective training. By leveraging just 2,048 NVIDIA H800 GPUs, DeepSeek-V3 achieves state-of-the-art performance. This achievement aligns with the commitment to advance AI through practical and scalable solutions, as previously demonstrated in the cost-effective architecture of Fire-Flyer AI-HPC (An et al., 2024). The practices and insights derived from DeepSeek-V3 demonstrate how existing hardware resources can be harnessed to their fullest potential, offering valuable lessons for the broader AI and HPC communities.
@@ 1.2. Objectives
This paper does not aim to reiterate the detailed architectural and algorithmic specifics of DeepSeek-V3, which are extensively documented in its technical report (DeepSeek-AI, 2024d). Instead, it adopts a dual perspective—spanning hardware architecture and model design—to explore the intricate interplay between them in achieving cost-efficient large-scale training and inference. By examining this synergy, we aim to provide actionable insights for scaling LLMs efficiently without sacrificing performance or accessibility.
@@ 1.3. Structure of this Paper
The remainder of this paper is organized as follows. Section 2 explores the design principles underpinning DeepSeek-V3 model architecture, highlighting key innovations such as Multi-head Latent Attention, Mixture-of-Experts optimizations and Multi-Token Prediction Module. Section 3 illustrates how our model architecture pursues low-precision computation and communication. Section 4 includes scale-up interconnection optimizations, discusses scale-up/scale-out convergence, and explores how hardware features influence parallelism and expert selection strategies. Section 5 focuses on scale-out network optimizations, including multi-plane network co-designs and low-latency interconnects. Besides current limitations and future suggestions mentioned in Section 3$\sim$5, Section 6 elaborates on more critical insights from DeepSeek-V3, and identifies directions for future hardware and model co-design.
@@ 2. Design Principles for DeepSeek Models
As shown in Figure 1, DeepSeek-V3 employs the DeepSeekMoE (DeepSeek-AI, 2024e) and Multi-head Latent Attention (MLA) (DeepSeek-AI, 2024c) architectures that have been proven effective in DeepSeek-V2 (DeepSeek-AI, 2024c). DeepSeekMoE unlocks the potential of MoE architecture, while MLA drastically reduces memory consumption by compressing Key-Value (KV) caches. In addition, DeepSeek-V3 incorporates FP8 mixed-precision training, significantly lowering computational costs and making large-scale training more practical without compromising model quality. To improve the inference speed, DeepSeek-V3 integrates speculative decoding based on its Multi-Token Prediction Module, which significantly increases the generation speed. Beyond model architecture, we also explored cost-efficient AI infrastructure by deploying a Multi-Plane two-layer Fat-Tree network to replace a traditional three-layer Fat-Tree topology, reducing cluster networking costs.
@@ 2.1. Memory Efficiency
LLMs generally require significant memory resources, with memory demands increasing by more than 1000% per year. In contrast, the growth rate of high-speed memory (e.g., HBM) capacity is much slower, typically less than 50% per year (Gholami et al., 2024). While multi-node parallelism is a viable solution to address memory limitations, optimizing memory usage at the source remains a crucial and effective strategy.
@@ 2.1.2. Reducing KV Cache with MLA
To address this bottleneck, we employ Multi-head Latent Attention (MLA) (DeepSeek-AI, 2024c) that compresses the KV representations of all attention heads into a smaller latent vector using a projection matrix, which is jointly trained with the model. During inference, only the latent vector needs to be cached, significantly reducing memory consumption compared to storing the KV cache for all attention heads.
@@
  • Shared KV (Grouped-Query Attention, GQA; Multi-Query Attention, MQA): Instead of maintaining separate KV pairs for each attention head, multiple heads share a single set of KV pairs, significantly compressing KV storage. Representative methods include GQA (Ainslie et al., 2023) and MQA (Shazeer, 2019).
  • Windowed KV: For long sequences, only a sliding window of KV pairs is retained in the cache, discarding results outside the window. While this reduces storage, it compromises long-context reasoning. Representative methods include Longformer (Beltagy et al., 2020) and related architectures.
  • Quantized Compression: KV pairs are stored using low-bit representations (Hooper et al., 2024; Liu et al., 2024; Kang et al., 2024), further reducing memory usage. Quantization achieves significant compression with minimal impact on model performance.
Table 1 compares the KV cache memory usage per token among DeepSeek-V3, Qwen-2.5 72B (Yang et al., 2024), and LLaMA-3.1 405B (AI@Meta, 2024b). By adopting MLA, DeepSeek-V3 achieves a significant reduction in KV cache size, requiring only 70 KB per token, substantially less than LLaMA-3.1 405B’s 516 KB and Qwen-2.5 72B’s 327 KB. This reduction highlights the efficiency of MLA in compressing KV representations compared to GQA-based methods. The ability to achieve such a significant reduction in memory consumption makes DeepSeek-V3 particularly well-suited for scenarios involving long-context processing and resource-constrained environments, enabling more scalable and cost-effective inference.
While reducing the size of the KV cache is a promising method for improving memory efficiency, the quadratic complexity inherent in Transformer-based autoregressive decoding remains a formidable challenge, especially for extremely long contexts. Recent research efforts, such as Mamba-2 (Dao and Gu, 2024) and Lightning Attention(Qin et al., 2024), investigate linear-time alternatives that offer new possibilities for balancing computational cost and model performance. In addition, approaches such as sparse attention (Yuan et al., 2025), which seek to compress and sparsely activate attention keys and values, represent another attempt at overcoming the computational challenges associated with attention. We look forward to collaborative progress with the broader community toward breakthroughs in this area.
2.2. Cost-Effectiveness of MoE Models
2.2.1. Reducing Computational Requirements for Training
@@ 2.2.2. Advantages for Personal Use and On-Premises Deployment
In a future where personalized LLM agents (Luo et al., 2025) become ubiquitous, MoE models offer unique advantages in single-request scenarios. Because only a subset of parameters is activated per request, memory and computational demands are greatly reduced. For example, DeepSeek-V2 (236B parameters) activates just 21B parameters during inference. This enables PCs with AI SoC chips (Apple, 2024; NVIDIA, 2025; AMD, 2025) to achieve nearly 20 tokens per second (TPS), or even twice that speed, which is more than sufficient for personal use. In contrast, dense models of similar capability (e.g., 70B parameters) typically reach only single-digit TPS on similar hardware.
Notably, the increasingly popular KTransformers (group and Approaching.AI, 2025) inference engine allows the complete DeepSeek-V3 model to run on a low-cost server equipped with a consumer GPU (costing approximately \$10,000), while still achieving nearly 20 TPS.
@@
2.3. Increasing Inference Speed
Inference speed encompasses both system-wide maximum throughput and single-request latency. To maximize throughput, our model is architected from the outset to leverage dual micro-batch overlap (DeepSeek-AI, 2025d; Zhao et al., 2025b), intentionally overlapping communication latency with computation. As demonstrated in our online inference system and supported by open-source profiling data (DeepSeek-AI, 2025d), we decouple the computation of MLA and MoE into two distinct stages. While one micro-batch executes a portion of MLA or MoE computation, the other micro-batch simultaneously performs the corresponding dispatch communication. Conversely, during the computation phase of the second micro-batch, the first micro-batch undergoes the combine communication step. This pipelined approach enables seamless overlap of all-to-all communication with ongoing computation, ensuring that the GPU remains fully utilized at all times. Moreover, in production, we adopt a prefill and decode disaggregation architecture (Zhong et al., 2024), assigning large batch size prefill and latency-sensitive decode requests to different expert parallelism group sizes. This strategy ultimately maximizes system throughput under real-world service conditions.
@@ 2.3.2. Inference Speed Limits
[…]
Here, dispatch uses FP8 (1 byte), while combine uses BF16 (2 bytes), and the hidden size of each token is approximately 7K. The factor 9 indicates that each token is transferred to 8 routed experts and 1 shared expert.
@@
Total Time Per $\text{Total Time Per Layer}=2\times 120.96\mu s=241.92\mu s$
@@
Total Inference $\text{Total Inference Time}=61\times 241.92\mu s=14.76\text{ms}$
@@
[…]
Assuming the computation time is equal to the communication time, this reduces the total inference time significantly, enabling a theoretical upper limit of over 0.82 ms TPOT, approximately 1200 tokens per second. While this figure is purely theoretical and has not been empirically validated, it vividly illustrates the transformative potential of high-bandwidth scale-up networks in accelerating large-scale model inference.
@@ 2.3.3. Multi-Token Prediction
Inspired by  Gloeckle et al. (2024), DeepSeek-V3 introduces a Multi-Token Prediction (MTP) framework, which simultaneously enhances model performance and improves inference speed. During inference, traditional autoregressive models generate one token at a decoding step, leading to sequential bottlenecks. MTP mitigates this issue by enabling the model to generate additional candidate tokens at a lower cost and verify them in parallel, similar to previous self-drafting-based speculative decoding approaches (Cai et al., 2024; Li et al., 2024). This framework significantly accelerates inference without compromising accuracy.
@@ 2.3.4. High Inference Speed for Reasoning Models and Test-Time Scaling
Test-time scaling in LLMs, exemplified by OpenAI’s o1/o3 series (OpenAI, 2024b, 2025), has enabled significant advances in mathematical reasoning, programming, and general reasoning by dynamically adjusting computational resources during inference. Subsequent models—including DeepSeek-R1 (DeepSeek-AI, 2025a), Claude-3.7 Sonnet (Anthropic, 2025), Gemini 2.5 Pro (Google, 2025), Seed1.5-Thinking (Seed, 2025), and Qwen3 (Team, 2025)—have adopted similar strategies and achieved notable improvements in these tasks.
For these reasoning models, high token output speed is of paramount importance. In reinforcement learning (RL) workflows—such as PPO (Schulman et al., 2017), DPO (Rafailov et al., 2024) and GRPO (Shao et al., 2024)—the necessity to rapidly generate large numbers of samples makes inference throughput a critical bottleneck. Likewise, prolonged reasoning sequences can increase user wait times, reducing the practical usability of such models. As a result, optimizing inference speed through synergistic hardware and software innovations is indispensable for advancing the efficiency of reasoning models. However, effective strategies for accelerating inference and expediting RL training remain active areas of investigation, as discussed in Section 2.1.3. We encourage the broader community to collaboratively explore and develop novel solutions to these ongoing challenges.

### v2

@@ 1.1. Background
Large Language Models (LLMs) have undergone rapid evolution in recent years, driven by iterative advancements in model design, computational power, and data availability. In 2024, groundbreaking models such as GPT4o (gpt4o), LLaMa-3 (llama3), Claude 3.5 Sonnet (claude35sonnet), Grok-2 (grok2), Qwen2.5 (qwen2.5), Gemini-2 (gemini2) and our DeepSeek-V3 (dsviii) have showcased remarkable progress, further narrowing the gap towards Artificial General Intelligence (AGI). As the Scaling Laws (scalinglaw) shows, increasing model size, training data, and computational resources leads to substantial improvements in model performance, underscoring the pivotal role of scaling in advancing AI capabilities. Collectively, these developments have ushered in an era where scaling model size and computational power is seen as the key to unlocking higher levels of intelligence.
@@
Recent developments, reasoning models such as OpenAI’s o1/o3 series models (openai_o1; openaio3), DeepSeek-R1 (dsr1), Claude-3.7 Sonnet (claude37sonnet), Gemini 2.5 Pro (gemini2_5), Seed1.5-Thinking (seed2025seed15thinkingadvancingsuperbreasoning) and Qwen3 (qwen3) have demonstrated not only the benefits conferred by large-scale architectures, but also the necessity of improving inference efficiency, particularly in handling longer contexts and achieving greater reasoning depth. These advancements underscore the need for faster and more efficient inference, consequently placing ever-increasing demands on computational resources.
@@
To meet these challenges, industry leaders such as Alibaba, ByteDance, Google, xAI and Meta have deployed colossal training clusters (10.1145/3579371.3589350; mudigere_software-hardware_2023-1; 10.1145/3651890.3672233; jiang_megascale_2024; 10.1145/3651890.3672265; xaicolossus), featuring tens or even hundreds of thousands of GPUs or TPUs. While such massive infrastructures have enabled the development of state-of-the-art models, their exorbitant costs present significant barriers for smaller research teams and organizations. Despite these barriers, open-source startups such as DeepSeek (dsvi; dsvii; dsviii; dscodervii; dsr1) and Mistral (mistral; mixtral8x22b) are also striving to develop state-of-the-art models. Among them, DeepSeek has especially demonstrated that effective software-hardware co-design can enable cost-efficient training of large models, leveling the playing field for smaller teams.
@@
Building on this tradition, DeepSeek-V3 (dsviii) represents a new milestone in cost-effective training. By leveraging just 2,048 NVIDIA H800 GPUs, DeepSeek-V3 achieves state-of-the-art performance. This achievement aligns with the commitment to advance AI through practical and scalable solutions, as previously demonstrated in the cost-effective architecture of Fire-Flyer AI-HPC (10793193). The practices and insights derived from DeepSeek-V3 demonstrate how existing hardware resources can be harnessed to their fullest potential, offering valuable lessons for the broader AI and HPC communities.
@@ 1.2. Objectives
This paper does not aim to reiterate the detailed architectural and algorithmic specifics of DeepSeek-V3, which are extensively documented in its technical report (dsviii). Instead, it adopts a dual perspective—spanning hardware architecture and model design—to explore the intricate interplay between them in achieving cost-efficient large-scale training and inference. By examining this synergy, we aim to provide actionable insights for scaling LLMs efficiently without sacrificing performance or accessibility.
@@ 1.3. Structure of this Paper
[…]
@@ 2. Design Principles for DeepSeek Models
As shown in Figure 1, DeepSeek-V3 employs the DeepSeekMoE (deepseekmoe) and Multi-head Latent Attention (MLA) (dsvii) architectures that have been proven effective in DeepSeek-V2 (dsvii). DeepSeekMoE unlocks the potential of MoE architecture, while MLA drastically reduces memory consumption by compressing Key-Value (KV) caches. In addition, DeepSeek-V3 incorporates FP8 mixed-precision training, significantly lowering computational costs and making large-scale training more practical without compromising model quality. To improve the inference speed, DeepSeek-V3 integrates speculative decoding based on its Multi-Token Prediction Module, which significantly increases the generation speed. Beyond model architecture, we also explored cost-efficient AI infrastructure by deploying a Multi-Plane two-layer Fat-Tree network to replace a traditional three-layer Fat-Tree topology, reducing cluster networking costs.
@@ 2.1. Memory Efficiency
LLMs generally require significant memory resources, with memory demands increasing by more than 1000% per year. In contrast, the growth rate of high-speed memory (e.g., HBM) capacity is much slower, typically less than 50% per year (10477550). While multi-node parallelism is a viable solution to address memory limitations, optimizing memory usage at the source remains a crucial and effective strategy.
@@ 2.1.2. Reducing KV Cache with MLA
To address this bottleneck, we employ Multi-head Latent Attention (MLA) (dsvii) that compresses the KV representations of all attention heads into a smaller latent vector using a projection matrix, which is jointly trained with the model. During inference, only the latent vector needs to be cached, significantly reducing memory consumption compared to storing the KV cache for all attention heads.
@@
  • Shared KV (Grouped-Query Attention, GQA; Multi-Query Attention, MQA): Instead of maintaining separate KV pairs for each attention head, multiple heads share a single set of KV pairs, significantly compressing KV storage. Representative methods include GQA (ainslie2023gqa) and MQA (mqa).
  • Windowed KV: For long sequences, only a sliding window of KV pairs is retained in the cache, discarding results outside the window. While this reduces storage, it compromises long-context reasoning. Representative methods include Longformer (Beltagy2020Longformer) and related architectures.
  • Quantized Compression: KV pairs are stored using low-bit representations (hooper2024kvquant; liu2024kivi; kang2024gear), further reducing memory usage. Quantization achieves significant compression with minimal impact on model performance.
@@
Table 1 compares the KV cache memory usage per token among DeepSeek-V3, Qwen-2.5 72B (qwen2.5), and LLaMA-3.1 405B (llama3_1_405b). By adopting MLA, DeepSeek-V3 achieves a significant reduction in KV cache size, requiring only 70 KB per token, substantially less than LLaMA-3.1 405B’s 516 KB and Qwen-2.5 72B’s 327 KB. This reduction highlights the efficiency of MLA in compressing KV representations compared to GQA-based methods. The ability to achieve such a significant reduction in memory consumption makes DeepSeek-V3 particularly well-suited for scenarios involving long-context processing and resource-constrained environments, enabling more scalable and cost-effective inference.
While reducing the size of the KV cache is a promising method for improving memory efficiency, the quadratic complexity inherent in Transformer-based autoregressive decoding remains a formidable challenge, especially for extremely long contexts. Recent research efforts, such as Mamba-2 (10.5555/3692070.3692469) and Lightning Attention(10.5555/3692070.3693758), investigate linear-time alternatives that offer new possibilities for balancing computational cost and model performance. In addition, approaches such as sparse attention (dsnsa), which seek to compress and sparsely activate attention keys and values, represent another attempt at overcoming the computational challenges associated with attention. We look forward to collaborative progress with the broader community toward breakthroughs in this area.
2.2. Cost-Effectiveness of MoE Models
2.2.1. Reducing Computational Requirements for Training
@@ 2.2.2. Advantages for Personal Use and On-Premises Deployment
In a future where personalized LLM agents (agentsurvey2025) become ubiquitous, MoE models offer unique advantages in single-request scenarios. Because only a subset of parameters is activated per request, memory and computational demands are greatly reduced. For example, DeepSeek-V2 (236B parameters) activates just 21B parameters during inference. This enables PCs with AI SoC chips (apple_m4; dgx_spark; amd_395) to achieve nearly 20 tokens per second (TPS), or even twice that speed, which is more than sufficient for personal use. In contrast, dense models of similar capability (e.g., 70B parameters) typically reach only single-digit TPS on similar hardware.
Notably, the increasingly popular KTransformers (ktransformers) inference engine allows the complete DeepSeek-V3 model to run on a low-cost server equipped with a consumer GPU (costing approximately \$10,000), while still achieving nearly 20 TPS.
@@
2.3. Increasing Inference Speed
Inference speed encompasses both system-wide maximum throughput and single-request latency. To maximize throughput, our model is architected from the outset to leverage dual micro-batch overlap (dsv3_profile_data; deepep2025), intentionally overlapping communication latency with computation. As demonstrated in our online inference system and supported by open-source profiling data (dsv3_profile_data), we decouple the computation of MLA and MoE into two distinct stages. While one micro-batch executes a portion of MLA or MoE computation, the other micro-batch simultaneously performs the corresponding dispatch communication. Conversely, during the computation phase of the second micro-batch, the first micro-batch undergoes the combine communication step. This pipelined approach enables seamless overlap of all-to-all communication with ongoing computation, ensuring that the GPU remains fully utilized at all times. Moreover, in production, we adopt a prefill and decode disaggregation architecture (298687), assigning large batch size prefill and latency-sensitive decode requests to different expert parallelism group sizes. This strategy ultimately maximizes system throughput under real-world service conditions.
@@ 2.3.2. Inference Speed Limits
$\text{Comm. Time}=(1\text{Byte}+2\text{Bytes})\times 32\times 9\times 7\text{K}/50\text{GB/s}=120.96\mu s$
Here, dispatch uses FP8 (1 byte), while combine uses BF16 (2 bytes), and the hidden size of each token is approximately 7K. The factor 9 indicates that each token is transferred to 8 routed experts and 1 shared expert. Network latency is not included in this calculation.
@@
$\text{Total Time Per Layer}=2\times 120.96\mu s=241.92\mu s$
@@
$\text{Total Inference Time}=61\times 241.92\mu s=14.76\text{ms}$
@@
$\text{Comm. Time}=(1\text{Byte}+2\text{Bytes})\times 32\times 9\times 7\text{K}/900\text{GB/s}=6.72\mu s$
Assuming perfect overlap between computation and communication, this would yield a theoretical upper limit of 0.82 ms TPOT, or approximately 1200 tokens per second. However, this figure is purely theoretical and does not account for the substantial drop in GPU efficiency at small batch sizes; in real deployments, actual throughput will be significantly lower. Nonetheless, this calculation vividly illustrates the transformative potential of high-bandwidth scale-up networks in accelerating large-scale model inference.
@@ 2.3.3. Multi-Token Prediction
Inspired by DBLP:conf/icml/GloeckleIRLS24, DeepSeek-V3 introduces a Multi-Token Prediction (MTP) framework, which simultaneously enhances model performance and improves inference speed. During inference, traditional autoregressive models generate one token at a decoding step, leading to sequential bottlenecks. MTP mitigates this issue by enabling the model to generate additional candidate tokens at a lower cost and verify them in parallel, similar to previous self-drafting-based speculative decoding approaches (DBLP:conf/icml/CaiLGPLCD24; DBLP:conf/icml/LiW0024). This framework significantly accelerates inference without compromising accuracy.
@@ 2.3.4. High Inference Speed for Reasoning Models and Test-Time Scaling
Test-time scaling in LLMs, exemplified by OpenAI’s o1/o3 series (openai_o1; openaio3), has enabled significant advances in mathematical reasoning, programming, and general reasoning by dynamically adjusting computational resources during inference. Subsequent models—including DeepSeek-R1 (dsr1), Claude-3.7 Sonnet (claude37sonnet), Gemini 2.5 Pro (gemini2_5), Seed1.5-Thinking (seed2025seed15thinkingadvancingsuperbreasoning), and Qwen3 (qwen3)—have adopted similar strategies and achieved notable improvements in these tasks.
For these reasoning models, high token output speed is of paramount importance. In reinforcement learning (RL) workflows—such as PPO (schulman2017proximalpolicyoptimizationalgorithms), DPO (rafailov2024directpreferenceoptimizationlanguage) and GRPO (shao2024deepseekmathpushinglimitsmathematical)—the necessity to rapidly generate large numbers of samples makes inference throughput a critical bottleneck. Likewise, prolonged reasoning sequences can increase user wait times, reducing the practical usability of such models. As a result, optimizing inference speed through synergistic hardware and software innovations is indispensable for advancing the efficiency of reasoning models. However, effective strategies for accelerating inference and expediting RL training remain active areas of investigation, as discussed in Section 2.1.3. We encourage the broader community to collaboratively explore and develop novel solutions to these ongoing challenges.